OpenStudy (anonymous):

Solve for x: 7/8 - 1/x = 3/4

OpenStudy (anonymous):

is it x=2?

OpenStudy (anonymous):

i got x = 8

OpenStudy (anonymous):

ok youre probably correct

OpenStudy (anonymous):

can u help with a few more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

solve for x: 2/5+3/5x=x+5/10

OpenStudy (anonymous):

confirm that the 5x is in the denominator

OpenStudy (anonymous):

i got x = –3 and x = 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 over 5x

OpenStudy (anonymous):

yes you are correct good job

OpenStudy (anonymous):

ok i have another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve for x: 2/x-2+7/x^2-4=5/x

OpenStudy (anonymous):

i got x=-4/3 and x=5

OpenStudy (anonymous):

have you learned imaginery numbers yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are u being serious

OpenStudy (anonymous):

yes. let me rework it

OpenStudy (anonymous):

ok my answer was probably wrong what did u get

OpenStudy (anonymous):